A fish is able to jump vertically out of the water with a speed of 4.9 m/s. How much time in seconds does it take for the fish to pass a point 0.5m above the water on the way up?
Answer #1

solution) u = 4.9 m/s h = o.sm t = ? - ut & h at a=-g h = ut - 18t2 to.s=0 - 4.9t 0.5 = (4.9) 11 (3.8) (+²) 0.5+ = (4.9) (t )